OpenStudy (anonymous):

What is the slope of the tangent line to the graph of h(t) at t=(1/3)π if h(t)=tan(t)?

OpenStudy (anonymous):

slope = derivative at that point! h '(t) = (sec(t))^2 @ t = (pi/3) --> h' (t) = 4 So the slope at that point is 4